Find the range of values of k for which the expression 3 - 4k - (k+3)x - x^2 will be negative for all real values of x.
[the answer is 3 < k < 7]

Answer by ikleyn(52928)

.
Find the range of values of k for which the expression 3 - 4k - (k+3)x - x^2 will be negative
for all real values of x.
[the answer is 3 < k < 7]
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        I just solved several similar problems yesterday and today, so at the regular studying process,
        it is just a time for a visitor to turn on his own mind, having several TEMPLATES from me.

The given quadratic function 3-4k - (k+3)x - x^2  has the leading coefficient -1 at x^2,
so it represents a downward parabola.


In order for this quadratic polynomial be negative at all real values of x, the necessary and 
sufficient condition is that the discriminant

    d = b^2 - 4ac 

be negative.  Then the square function has no real roots and remains negative for all real values of x.


So, we write the discriminant

    d = (-(k+3))^2 - 4*(-1)*(3-4k) = (k+3)^2 + 4*(3-4k) = k^2 + 6k + 9 + 12 - 16k = k^2 - 10k + 21.


It can be factored

    d = (k-7)*(k-3).


So, the discriminant is negative

    (k-7)*(k-3) < 0


if and only if  the parameter "k" is between the roots  3 < k < 7,
when the factor (k-7) is negative, while the factor (k-3) is positive.


At this point, the problem is fully solved.


ANSWER.  The range of values of k for which the expression  3 - 4k - (k+3)x - x^2  will be negative 

         for all real values of x  is  3 < k < 7.

Solved.